HIGH PRECISION MID-IR SPECTROSCOPY OF ^{12}C^{16}O_{2} NEAR 4.3 μm

\underline{WEI-JO TING}, PEI-LING LUO,CHIEH-HSING CHUNG, Dept. of Physics, National Tsing Hua University, Hsinchu, Taiwan 30013, R.O.C; HSHAN-CHEN CHEN, Inst. of Photonics Technologies, National Tsing Hua University, Hsinchu, Taiwan 30013, R.O.C; YU-HUNG LIEN, and JOW-TSONG SHY, Dept. of Physics, National Tsing Hua University, Hsinchu, Taiwan 30013, R.O.C.

We have observed the sub-Doppler saturation spectrum of the ^{12}C^{16}O_{2} near 4.3 μm using a mW-level DFG (Difference Frequency Generation) source. The DFG radiation is generated by a 1-W Ti:sapphire laser and a Nd:YAG laser amplified by 10-W fiber amplifier in a 50-mm long PPLN (Periodically Poled Lithium Niobate) crystal. We are able to generate 2 mW DFG power at 4.3 μm. The Nd:YAG laser is frequency-doubled, and frequency stabilized on one ^{127}I_{2} hyperfine transition. The Ti:sapphire laser is locked onto the center of CO_{2} transition and its frequency is measured by an OFC (Optical Frequency Comb). In this talk, we will report our recent measurements on the high J 0001 → 0000 fundamental band transitions and the hot band 0111 → 0110 transitions. To increase the signal-to-noise ratio of the observed spectra, the CO_{2} absorption cell in heated to a temperature > 500^{∘}C.